wolfmaster:

Nadia is mountain climbing. She started at an altitude of 19.26 feet below sea level and then changed her altitude by climbing a total of 5,437.8 feet up from her initial position. What was Nadia’s altitude at the end of her climb? 3,511.8 feet above sea level 5,245.2 feet above sea level 5,418.54 feet above sea level 5,457.06 feet above sea level

Imagine:

\[(-19.26+x= 5,437.8)\]

Imagine:

ꜱᴏʟᴠᴇ ꜰᴏʀ "x"

wolfmaster:

Okay thanks

Imagine:

Anytime!

Imagine:

Do you know what to do from here?

Imagine:

\[(-19.26+3,511.8)=3492.54\]

Imagine:

\[(-19.26+5,245.2)=5225.94\]

Imagine:

\[(-19.26+5,418.54)=5399.28\]

Imagine:

\[(-19.26+5,457.06)=5437.8\]